Find all solutions of the given differential equations and then find the particular solutions for which a point (x,y) is given:


dy/dx = x


1
Expert's answer
2021-09-02T00:34:26-0400

We may rewrite the equation in form

"dy = x\cdot dx,"

next, we integrate both parts and get

"\int dy = \int x\cdot dx, \\
y = \dfrac12\cdot x^2 + c."


If the point "(x_0, y_0)" is given, we may obtain the constant c:

"c = y_0 - \dfrac12\cdot x_0^2\,,"

so the solution will be

"y = \dfrac12x^2 + y_0 - \dfrac12x_0^2\,."